Verify whether f(x) = (8/7) (1/2)^x



, x = 1,2,3 is a probability mass function. If it is,




determine the following probabilities.



(a) P(X ≤ 1)



(b)P(X > 1)



(c) P(2 < X < 6)



(d)P(X ≤ 1 or X > 1)

1
Expert's answer
2022-03-08T01:04:18-0500
"f(x)=\dfrac{8}{7}(\dfrac{1}{2})^x,x=1,2,3"

"\dfrac{8}{7}(\dfrac{1}{2})^x>0,x=1,2,3"


"\dfrac{8}{7}(\dfrac{1}{2})^1+\dfrac{8}{7}(\dfrac{1}{2})^2+\dfrac{8}{7}(\dfrac{1}{2})^3=1"




"P(X=1)=\dfrac{8}{7}(\dfrac{1}{2})"

"P(X=2)=\dfrac{8}{7}(\dfrac{1}{4})"

"P(X=3)=\dfrac{8}{7}(\dfrac{1}{8})"

"f(x)" is a probability mass function.


(a)


"P(X\le1)=\dfrac{4}{7}"

(b)


"P(X>1)=1-\dfrac{4}{7}=\dfrac{3}{7}"

(c)


"P(2<X<6)=P(X=3)=\dfrac{1}{7}"

(d)


"P(X\le 1\ or\ X>1)=1"